OpenStudy (anonymous):

boiling point of water with 3 tablespoons of salt MEDAL & FAN

OpenStudy (aaronq):

you want the exact BP of water after adding 3 tbsp of "table salt" (as in NaCl)? use the BP elevation equation (and the constant Kb) \(\Delta T=i*m*K_b\) where \(\Delta T\) is the change in temp from the pure solvent's BP \(i\) is the vant hoff constant (for NaCl it is equal to 2\) \(m\) is the molality of the solution \(\sf m=\dfrac{moles~of~solute}{Liters~of~solution}\)

OpenStudy (anonymous):

NaCl

OpenStudy (aaronq):

k cool. now you need to do a few steps. First convert the units of salt to mass. next convert the mass of NaCl to moles of NaCl. \(\sf moles =\dfrac{mass}{Molar~mass}\) next find the molality (see first post) next find the change in temp and add it to the original BP, that is: \(\sf New~BP=\Delta T+100~^oC\)